# CafkNet: GNN-Empowered Forward Kinematic Modeling for Cable-Driven Parallel Robots

### 28 Feb 2024 v1

@@ VI-C2 Real Experimental Data
To evaluate the sim2real capability of CafkNet, we design three methods based on the training data, as presented in Tab. V. Furthermore, in the testing, cable lengths from either the clean data or noise data are inputted into the trained model. Then it solves for the end-effector poses and compares them with the corresponding poses from the clean data and noise data to compute the average error, respectively.
@@
  • sim&real2real: This method introduces composite training data, which consists of simulated data and $80\%$ of clean data. The testing data remains consistent with the real2real method.
The resulting trajectories solved by CafkNet under these three methods are presented in Fig. 6, and their accuracy results are given in Tab. V. The experimental findings demonstrate that CafkNet, trained using composite data consisting of simulated data and $80\%$ of the clean data, yields end-effector poses that closely align with real machine data. This indicates its effectiveness in bridging the sim2real gap in practical applications of CDPRs. Specifically, when testing on clean data, as expected, the sim2real method exhibits the lowest error, followed by sim&real2real, while real2real performs the poorest. This can be attributed to the fact that the quantity of simulated data ($10000$) is ten times that of the clean data ($1000$). Consequently, the sim2real method has more information about end-effector poses and cable lengths, which contributes to its superior performance on clean data (error in $1.7$ mm). Compared to the real2real method, the sim&real2real method supplements the data, resulting in a decrease in modeling error from $2.7$ to $2.4$ mm.
When directly transferring the model to unseen noise data, the sim&real2real method achieves the best performance, with an accuracy error as low as $1.3$ mm. This can be attributed to the method’s rich dataset, which not only includes a large range of motion information but also incorporates specific information about the target trajectory. These factors contribute to the model’s ability to minimize FK-solving errors during sim2real transfer.
@@
Figure 4: Modeling accuracy of CafkNet working on artificial noise data. Here errors in case 1 to case 3 are generated by $\epsilon\sim N(0,5^{2})$, and errors in case 4 to case 6 are generated by $\epsilon\sim N(0,10^{2})$. $N$ refers to the Gaussian distribution.
Figure 5: Data sources of Sim2Real experiment for the configuration of ExpC4. (i) Simulated data. In the simulator, there are $100$ random trajectories and $100$ samplings per trajectory. (ii) Clean data. In the simulator, we have devised a target trajectory with the objective of controlling the real CDPR to execute it. Here $1000$ samplings are generated. (iii) Noise data. In the real setup, the end-effector of CDPR follows the target trajectory. During the experiment, a total of 1103 data are collected. Above all, each data composes the end-effector pose ($x,y,\theta$) and the corresponding length of $4$ cables.
TABLE V: Sim2Real Experiment Details and Results. Cases i, ii, iii are revealed in Fig. 5. Bold: minimum error in each testing data.
@@
sim&real2real $\text{i}+80\%$ ii $20\%$ ii iii 2.4 1.3
@@ VII Conclusion
This paper presents CafkNet, a GNN-empowered FK solver for CDPRs. It is the first work to utilize GNNs for solving the FK problem in CDPRs. The study begins by introducing a graph structure that represents the topology of CDPRs. Subsequently, GNNs are employed to solve the FK problem by learning pose-length data. The experimental results demonstrate that CafkNet can learn the inherent topological relationships of CDPRs and accurately solve the FK issue. The results are validated across different types of CDPRs, including spatial CDPRs and planar CDPRs, as well as under various configurations such as under-constrained, fully-constrained, and over-constrained CDPRs. The validation is conducted in both simulation environments and real-world scenarios as well.

### 05 Mar 2024 v2

@@ VI-C2 Real Experimental Data
To evaluate the sim2real capability of CafkNet, we design three methods based on the training data, as summarized in Tab. IV. Then, in the testing, cable lengths from either the clean data or noise data are inputted into the trained model. Then it solves for the end-effector poses and compares them with the corresponding poses from the clean data and noise data to compute the average error, respectively. Three methods are described in detail below.
@@
  • $sim\&real2real$: This method introduces training data including all simulated data and $80\%$ of clean data. The testing data remains consistent with the $real2real$ method.
The resulting trajectories solved by CafkNet under these three methods are presented in Fig. 6, and their accuracy results are given in Tab. IV. The experimental findings demonstrate that CafkNet, trained using composite data consisting of simulated data and $80\%$ of the clean data, yields end-effector poses that closely align with real machine data. This indicates its effectiveness in bridging the sim2real gap in practical applications of CDPRs. Specifically, when testing on clean data, as expected, the ${sim2real}$ method exhibits the lowest error, followed by ${sim\&real2real}$, while ${real2real}$ performs the poorest. This can be attributed to the fact that the quantity of simulated data ($10000$) is ten times that of the clean data ($1000$). Consequently, the ${sim2real}$ method has more information about end-effector poses and cable lengths, which contributes to its superior performance on clean data (error in $1.7$ mm). Compared to the ${real2real}$ method, the ${sim\&real2real}$ method supplements the data, resulting in a decrease in modeling error from $2.7$ to $2.4$ mm.
When directly transferring the model to unseen noise data, the method ${sim\&real2real}$ achieves the best performance, with an accuracy error as low as $1.3$ mm. This can be attributed to the method’s rich dataset, which not only includes a large range of motion information but also incorporates specific information about the target trajectory. These factors contribute to the model’s ability to minimize FK-solving errors during sim2real transfer.
@@
TABLE IV: Sim2Real Experiment Details and Results. Cases i, ii, iii are revealed in Fig. 5. Bold: minimum error in each testing data.
@@
$sim\&real2real$ $\text{i}+80\%$ ii $20\%$ ii iii 2.4 1.3
@@ VII Conclusion
This paper presents the CafkNet, a GNN-empowered FK solver for CDPRs. It is the first work to utilize GNN for FK problem-solving in CDPRs. The study begins by introducing a graph structure that represents the topology of CDPRs, followed by network learning on the graph via message propagation block. The experimental results demonstrate that CafkNet can learn the inherent topological relationships of CDPRs and efficiently solve the FK issue. The results are validated across different types of CDPRs, including spatial CDPRs and planar CDPRs, as well as under configurations such as under-constrained, fully-constrained, and over-constrained CDPRs. The validation is conducted in both simulation environments and real-world scenarios as well.